# RWZC: A Model-Driven Approach for Learning-based Robust Wyner-Ziv Coding

### v1

@@ III-D Warping-predicted reconstruction with SSI
For the high-quality reconstruction, a direct idea is to blend masked image $\tilde{\bm{x}}$ and a warped side image $\bm{y}$. One can find the above modules present a transmitted image portion at the receiver, while the complement portion can also be obtained by the correlated side image $\bm{y}$. However, the pixel-wise blending will inevitably result in two problems. The first problem has been illustrated in the second row in Fig. 5 in Sec. III-B, since the varying $\bm{h}$ at the receiver influences the blending performance owing to a distorted perspective transform and an inaccurate SSI. Second, unlike the pure compression framework, transmission inevitably incurs noise which affects the quality of the received image, and further influences the final combination of $\tilde{\bm{x}}$ and $\bm{y}$. Specifically, due to the resolution differences and noise interference at the masked edges, distortion will occur while ultimately resulting in discontinuities in the reconstructed image, especially for extreme low rate and large noisy communication. These two problems though impacts the perceptual performance slightly, while it is crucial for some objective metrics.
@@
where $\mathcal{L}_{s}$ denotes the boundary and smoothness loss [31] for seamless blending, while $b$ is the parameter to trade-off these two terms. Fig. 9 shows the effect of the learning-based $\mathcal{W}_{\eta}$, in which the first row completes the blending with a direct perspective transform according to $\bm{h}$. One can easily find the obvious pixel deviation at the stitching position caused by resolution difference of two portions. This problem can be solved by the predicted compensation and blur operation around the seam via our network.
@@ III-E Loss function and training algorithm
In this work, our proposed framework has three main learning-based neural networks, namely adaptive masking $\mathcal{M}_{\eta}$, JSC codec $F_{\theta}$ and $F_{\xi}$, and warping prediction reconstruction $\mathcal{W}_{\zeta}$. For the training process, we jointly train the three networks for an optimal performance, in which the combination of PTL-based mask and variational autoencoder (VAE)-based JSC codec ensures the performance against varying homography and channel interference, while prediction-based warping completes a pixel-level refinement. Moreover, we remark that the joint training enjoys a better performance despite a higher temporal complexity than the separated training method, while it is still lower than its CNN-based or transformer-based competitors, due to the linear-layer PTL and point-to-point simple autoencoder design. The loss functions can be formulated as follows.
@@ IV Experiments
In this section, performance of our proposed framework will be verified. We select two state-of-the-art competitors, which consider efficient compression and channel-robust transmission at a WZ coding scenario. In conclusion, for the reconstruction task of $\bm{x}$, our framework outperforms the baselines on perceptual metrics significantly, while maintaining comparable performance on objective metrics, especially for extremely low rate. Moreover, comparison of complexity is presented to demonstrate the advantage of the light-weight codec design in our work.
@@ IV-A2 Evaluation Metrics
To evaluate the performance of the decoder-only side information empowered framework, we adopt the widely-used metric PSNR, which is used to assess the quality of reconstructed images compared to the original image. Except for the point-wised object metric, we also used MS-SSIM and learned perceptual image patch similarity (LPIPS) for evaluation, which are both used as the proxy for the human perception. Among them, MS-SSIM is calculated with luminance, contrast, and structure between two images, while LPIPS computes the distance in the feature space based on a pretrained network.
@@ IV-A3 Model and Training Details
We model the complex AWGN channel with noise sampled from $\mathcal{CN}(\bm{0},\sigma_{n}^{2}\bm{I}_{k})$, normalize the transmitting power and control $\sigma_{n}^{2}$ for SNR computation. In RWZC framework, we adopt algorithms SURF and RANSAC 5.0 for feature matching and defects exclusion in $\mathcal{H}$, respectively. For mask generating strategy $\mathcal{M}_{\eta}$, we choose bicubic interpolation for PIL and the list size of trainable homography $\bm{h}^{*}$ equals to 1. For JSCC codec $F_{\theta}$ and $F^{-1}_{\xi}$, input/output/kernal details are labeled in Fig. 7, and we construct a set of 16 different FC layers, whose output size ranges from 4 to 128 to realize the transmission rate adaption. For final stitching in $\mathcal{W}_{\zeta}$, Poisson blending is used for seamless reconstruction. Moreover, we resize the images in 370$\times$740 and 256$\times$256 for KITTI Stereo and UDIS-D datasets, respectively. Besides, for the training, $i_{\mathrm{max}}$ equals to 400, $\lambda_{1}=0.7$ and $\lambda_{2}=0.1$ while the learning rate is set as $1e^{-4}$. The framework is optimized with Adam [35] and the distortion measure is selected as MS-SSIM.
@@ IV-A4 Baselines
Mital+capacity denotes the scheme of NDIC combined with capacity achieving code. It is a separation-based framework, in which the compression is based on the entropy model from [36] with a common feature extraction at decoder, and the transmission is assumed error-free at the maximun achievable rate, computed by $\frac{k}{n}$ times the channel capacity[37]. Specifically, the method of Mital et al. utilizes learned image compression with hyperprior as the backbone and loss function characterizing the trade-off between rate and distortion. It should be noticed that, in Mital+capacity, the CBR cannot be less than $0.027$ since a too small trade-off parameters leads to the non-convergence of training process.
IV-B Evaluation on Stationary Parallax Dataset
@@
Fig. 10 presents the reconstruct performance of all frameworks over different bandwidth usages on the dataset with stationary parallax. In general, our method performs as good as DeepJSCC-WZ and outperforms Mital+capacity, in terms of PSNR and SSIM metrics. Specifically, the PSNR of RWZC can reach 25.5 dB when CBR is around 0.03 (at 3 dB channel SNR), while DeepJSCC-WZ has a slightly weaker performance of 25 dB and Mital+capacity only obtains 22.1 dB. The case turns a bit different when CBR climbs to 0.06, while DeepJSCC-WZ overtakes our scheme around 0.02 dB. Moreover, the MS-SSIM of RWZC achieves 0.956 while DeepJSCC-WZ has 0.952 and Mital+capacity has only 0.918, when CBR approaches 0.03. Besides, our method shows an overwhelming trends compared with DeepJSCC-WZ in LPIPS metric, which appears an average 0.08 improvement as shown in the third plot.
@@
In Fig. 12, visual results are presented to show the specific performance. The stationary homography is illustrated in the third column (marked as Masking(SSI)), where the $\bm{h}$-generated mask remains almost invariant for most image pairs. The visualization also validates the comparable performance of RWZC with DeepJSCC-WZ. One can conclude that for the KITTI stereo dataset, our proposed method does not obtain a significant advantage when compared to existing methods, (even negative gain on PSNR/MS-SSIM compared with DeepJSCC-WZ) except for the perceptual metric, no matter against CBR or SNR. The reason lies on that the adequate information carried by the side information in the stationary parallax scenario, which offers effective improvement for data-driven-based feature extraction. However, we argue that it is not the same case when the parallax is not stationary, which will be shown in the next section.
@@ IV-C Evaluation on Parallax-Varying Dataset
Figure 16: Cross-Dataset testing of proposed method with the baselines by fixing CBR0.031absent0.031$\approx$ 0.031≈ 0.031, where cross-dataset testing(abbreviated by c.t.) means that the framework is trained on the KITTI stereo and tested on the UDIS-D.
@@ IV-D Cross-dataset Testing
Moreover, we conduct cross-dataset testing across all baselines. This approach is commonly used in machine learning to evaluate a model’s generalization ability, ensuring it performs well on unseen data from potentially different domains, distributions, or environments. Specifically, we fix the CBR around 0.031, train RWZC and the other baselines on the KITTI stereo dataset, and test them on the UDIS-D dataset, where the images are resized to 370$\times$740.
Figure 16 presents the performance comparison among different baselines under cross-dataset testing. We observe that our framework outperforms competitors in terms of all metrics—PSNR, MS-SSIM, and LPIPS—when faced with image samples from unfamiliar distributions. Additionally, it is important to note that cross-dataset testing inevitably leads to some performance degradation. Therefore, we provide a comparison between the in-dataset testing (where both training and testing are done on the same dataset) and cross-dataset testing for each baseline. Moreover, the percentage above each bar indicates the performance degradation. For example, considering the PSNR metric with SNR fixed at 1dB (as shown in the left plot of Figure 16), RWZC with cross-dataset testing achieves 26.59 dB, while the original framework achieves 26.952 dB, representing a $1.3\%$ degradation. Meanwhile, DeepJSCC-WZ shows a degradation of about $6.5\%$, and Mital+Capacity exhibits a $4.4\%$ degradation. We conclude that a smaller percentage indicates less performance degradation, reflecting a more robust framework when facing unfamiliar data distributions.
@@ IV-E Ablation studies
The comparisons among these schemes in terms of MS-SSIM and LPIPS over SNR are shown in Fig. 17. The plots show that the learnable masking strategy contributes the most to the final performance, while the rate-adaptive transmission follows, and the warping prediction reconstruction contributes the least, since the generalization to unfamiliar SSI is essential to the framework. The trends turns a slight different in LPIPS comparison, where the rate adaption and warping prediction modules hardly influences the performance since the shifting of a few pixels will not affect the human perception.
@@ V Conclusion
In this work, we propose a novel learning-based Wyner-Ziv coding framework that is robust to non-stationary source correlations. Unlike existing data-driven approaches, our framework establishes an affine model between images, effectively separating the overlapping information of the transmitted image from the side image, enabling reliable and efficient communication. These schemes have been shown to outperform state-of-the-art competitors, particularly in terms of perceptual metrics.
For future work, we plan to explore different modeling approaches for source correlation in the RWZC framework, which may involve non-linear correlations, such as those arising from illuminations and depth, rather than just affine correlation. Non-linear correlation modeling better aligns with real-world signal sources but will require more advanced network architecture design. Additionally, we aim to extend our framework to better accommodate varying channel environments, including fading channels and MIMO systems.

### v2

@@ III-D Warping-predicted reconstruction with SSI
For the high-quality reconstruction, a direct idea is to blend masked image $\tilde{\bm{x}}$ and a warped side image $\bm{y}$. One can find the above modules present a transmitted image portion at the receiver, while the complement portion can also be obtained by the correlated side image $\bm{y}$. However, the pixel-wise blending will inevitably result in two problems. The first problem has been illustrated in the second row in Fig. 5 in Sec. III-B, since the varying $\bm{h}$ at the receiver influences the blending performance owing to a distorted perspective transform and an inaccurate SSI. Second, unlike the pure compression framework, transmission inevitably incurs noise which affects the quality of the received image, and further influences the final combination of $\tilde{\bm{x}}$ and $\bm{y}$. Specifically, due to the resolution differences and noise interference at the masked edges, distortion will occur while ultimately resulting in discontinuities in the reconstructed image, especially for communication environments with extremely low rates and strong interference. These two problems though impacts the perceptual performance slightly, while it is crucial for some objective metrics.
@@
where $\mathcal{L}_{s}$ denotes the boundary and smoothness loss [31] for seamless blending, while $b$ is the parameter to trade-off these two terms. Fig. 9 shows the effect of the learning-based $\mathcal{W}_{\eta}$, in which the first row completes the deirect blending with a perspective transformed side image according to $\bm{h}$. One can easily find the obvious pixel deviation at the stitching position caused by resolution difference of two portions. In general, this problem can be solved by the predicted compensation and blur operation around the seam via our network.
@@ III-E Loss function and training algorithm
In this work, our proposed framework has three main learning-based neural networks, namely adaptive masking $\mathcal{M}_{\eta}$, JSC codec $F_{\theta}$ and $F_{\xi}$, and warping prediction reconstruction $\mathcal{W}_{\zeta}$. For the training process, we jointly train the three networks for an optimal performance, in which the combination of PTL-based mask and variational autoencoder (VAE)-based JSC codec ensures the performance against varying homography and channel interference, while prediction-based warping completes a pixel-level refinement. Moreover, we remark that the joint training though suffers from a higher complexity with its separated counterpart, it always enjoys a better performance. In specific, the loss functions can be formulated as follows.
@@ IV Experiments
In this section, performance of our proposed framework will be verified. We select two state-of-the-art competitors, which consider efficient compression and channel-robust transmission at a WZ coding scenario, respectively. In conclusion, for the reconstruction task of $\bm{x}$, our framework outperforms the baselines on perceptual metrics significantly, while maintaining comparable performance on objective metrics, especially for extremely low rate. Moreover, comparison of complexity is presented to demonstrate the advantage of the light-weight codec design in our work.
@@ IV-A2 Evaluation Metrics
To evaluate the performance of the decoder-only side information empowered framework, we adopt the widely-used metric PSNR, which is used to evaluate the quality of reconstructed images compared to the original image. Except for the point-wised object metric, we also used MS-SSIM and learned perceptual image patch similarity (LPIPS) for evaluation, which are both used as the proxy for the human perception. Among them, MS-SSIM is calculated with luminance, contrast, and structure between two images, while LPIPS computes the distance in the feature space based on a pretrained network.
@@ IV-A3 Model and Training Details
We model the complex AWGN channel with noise sampled from $\mathcal{CN}(\bm{0},\sigma_{n}^{2}\bm{I}_{k})$, normalize the transmitting power and control $\sigma_{n}^{2}$ for SNR computation. In RWZC framework, we adopt algorithms SURF and RANSAC 5.0 for feature matching and defects exclusion in $\mathcal{H}$, respectively. For mask generating strategy $\mathcal{M}_{\eta}$, we choose bicubic interpolation for PIL and the list size of trainable homography $\bm{h}^{*}$ equals to 1. For JSCC codec $F_{\theta}$ and $F^{-1}_{\xi}$, input/output/kernal details are labeled in Fig. 7, and we construct a set of 16 different FC layers, whose output size ranges from 4 to 128 to realize the transmission rate adaption. For final stitching in $\mathcal{W}_{\zeta}$, Poisson blending is used for seamless reconstruction. Moreover, we resize the images in 370$\times$740 and 256$\times$256 for KITTI Stereo and UDIS-D datasets, respectively. Besides, for the training, $i_{\mathrm{max}}$ equals to 400, $\lambda_{1}=0.7$ and $\lambda_{2}=0.1$ while the learning rate is set as $1e^{-4}$. The framework is optimized with Adam [35] and trained with the distortion measure of MS-SSIM.
@@ IV-A4 Baselines
Mital+capacity denotes the scheme of NDIC combined with capacity achieving code. It is a separation-based framework, in which the compression is based on the entropy model from [36] with a common feature extraction at decoder, and the transmission is assumed error-free at the maximun achievable rate, computed by $\frac{k}{n}$ times the channel capacity[37]. Specifically, the method of Mital et al. utilizes learned image compression with hyperprior as the backbone and loss function characterizing the trade-off between rate and distortion. It should be noticed that, in Mital+capacity, the CBR cannot be less than $0.027$ since too strict compression may lead to non-convergence during the training process.
@@ IV-B Evaluation on Stationary Parallax Dataset
Fig. 10 presents the reconstruction performance of all frameworks over different bandwidth usages on the dataset with stationary parallax. In general, our method performs as good as DeepJSCC-WZ and outperforms Mital+capacity, in terms of PSNR and MS-SSIM metrics. Specifically, the PSNR of RWZC can reach 25.5 dB when CBR is around 0.03 (at 3 dB channel SNR), while DeepJSCC-WZ has a slightly weaker performance of 25 dB and Mital+capacity only obtains 22.1 dB. The case turns a bit different when CBR climbs to 0.06, while DeepJSCC-WZ overtakes our scheme around 0.02 dB. Moreover, the MS-SSIM of RWZC achieves 0.956 while DeepJSCC-WZ has 0.952 and Mital+capacity has only 0.918, when CBR approaches 0.03. Besides, our method shows an overwhelming trends compared with DeepJSCC-WZ in LPIPS metric, which appears an average 0.08 improvement as shown in the third plot.
@@
In Fig. 12, visual results are presented to show the specific performance. The stationary homography is illustrated in the third column (marked as Masking(SSI)), where the $\bm{h}$-generated mask remains almost invariant for most image pairs. The visualization also validates the comparable performance of RWZC with DeepJSCC-WZ. One can conclude that for the KITTI stereo dataset, our proposed method does not obtain a significant advantage when compared to existing methods, (even negative gain on PSNR/MS-SSIM compared with DeepJSCC-WZ) except for the perceptual metric, no matter against CBR or SNR. The reason lies on that the adequate information carried by the side information in the stationary parallax scenario, which offers effective improvement for data-driven-based feature extraction. However, we want to clarify that it is not the same case when the parallax is not stationary, which will be shown in the next section.
@@ IV-C Evaluation on Parallax-Varying Dataset
Figure 16: Cross-Dataset test of proposed method with the baselines by fixing CBR0.031absent0.031$\approx$ 0.031≈ 0.031, where herein this test means that the framework is trained on the KITTI stereo and tested on the UDIS-D.
@@ IV-D Cross-dataset Testing
Moreover, we conduct cross-dataset testing across all baselines. This approach is commonly used in machine learning to evaluate a model’s generalization ability, ensuring it performs well on unseen data from potentially different domains, distributions, or environments. Specifically, we fix the CBR around 0.031, train RWZC and the other baselines on the KITTI stereo dataset, and test them on the UDIS-D dataset, where the images are resized to both 370$\times$740.
Figure 16 presents the performance comparison among different baselines under cross-dataset testing. We observe that our framework outperforms competitors in terms of all metrics—PSNR, MS-SSIM, and LPIPS—when faced with image samples from unfamiliar distributions. Additionally, it is important to note that cross-dataset testing inevitably leads to some performance degradation. Therefore, we provide a comparison between the in-dataset testing (where both training and testing are done on the same dataset) and cross-dataset testing for each baseline. Moreover, the percentage above each bar indicates the performance degradation. For example, considering the PSNR metric with SNR fixed at 1dB (as shown in the left plot of Figure 16), RWZC with cross-dataset testing achieves 26.59 dB, while the original framework achieves 26.952 dB, representing a $1.3\%$ degradation. Meanwhile, DeepJSCC-WZ shows a degradation of about $6.5\%$, and Mital+Capacity exhibits a $4.4\%$ degradation. We conclude that a smaller percentage indicates less performance degradation, reflecting a stronger robustness when facing unfamiliar data distributions of our framework.
@@ IV-E Ablation studies
The comparisons among these schemes in terms of MS-SSIM and LPIPS over SNR are shown in Fig. 17. The plots show that the learnable masking strategy contributes the most to the final performance, while the rate-adaptive transmission follows, and the warping prediction reconstruction contributes the least, since the generalization to unfamiliar SSI is essential to the framework. The trends turn a slight different in LPIPS comparison, where the warping prediction module hardly influences the performance since the shifting of a few pixels will not affect the human perception.
@@ V Conclusion
In this work, we propose a novel learning-based Wyner-Ziv coding framework that is robust to non-stationarily correlated sources. Unlike existing data-driven approaches, our framework establishes an affine model between images, effectively separating the overlapping information of the transmitted image from the side image, enabling reliable and efficient communication. These schemes have been shown to outperform state-of-the-art competitors, particularly in terms of perceptual metrics.
For future work, we plan to explore different modeling approaches for source correlation in the RWZC framework, which may involve non-linear correlations, such as those arising from illuminations and depth, rather than just affine correlation. Non-linear correlation modeling can better align with real-world sources but require more advanced network architecture design. Additionally, we aim to extend our framework to better accommodate varying channel environments, including fading channels and MIMO systems.